suppose electrical energy cost 25 cents per kilo watt hour. Calculate the cost of the electrical service to a house which uses four 40w hours and three 60w lamps for 100 hours
1
Expert's answer
2020-01-27T10:01:07-0500

Solution. Find the total lamp power


"P=4\times 40+3\times 60=160+180=340W"

Find the amount of energy consumed in kilowatt hours


"E=\frac {340}{1000}kW \times 100h=34 kW h"

According to the condition of the problem energy cost 25 cents per kilo watt hour. Therefore the cost of the electrical service to a house is equal to


"cost = 34kWh \times 0.25 \frac {dollars} {kWh}=8.5 dollars"

Answer. 8.5 $.